OpenStudy (anonymous):

What is the difference between Reflexive, Symmetric, Transitive, Additive, and Multiplicative Axiom?

OpenStudy (anonymous):

think for reflexive of a mirror symmetric can it fold on both sides transitive can you move it

OpenStudy (anonymous):

Reflexive Axiom is a number equal to itself. (a=a) Symmetric Axiom are symmetric around the equal sign. If a=b, then b=a. Transitive Axiom is if a=b and b=c, then a=c. Additive Axiom is if a=b and c=d, then a+c=b+d Multiplicative Axiom is if a=b and c=d, and ac=bd. The multiplicative axiom follows the additive axiom.